OpenStudy (anonymous):

Find the determinant of the following matrix 2 -3 4 5 1 -2 10 3 -1

OpenStudy (anonymous):

\[\left| \det \right| = 109\]

OpenStudy (anonymous):

2(-1+6) +3(-5 + 20) + 4(15 - 10) = 10 + 45 + 20 = 75

OpenStudy (anonymous):

\[\left|\begin{matrix} 2&-3&4\\5&1&-2\\10&3&-1 \end{matrix}\right|=2\left|\begin{matrix} 1&-2\\3&-1 \end{matrix}\right|+3\left|\begin{matrix} 5&-2\\10&-1 \end{matrix}\right|+4\left|\begin{matrix} 5&1\\10&3 \end{matrix}\right|\]

OpenStudy (anonymous):

i dont get how you get the matrix for 3 and 4

OpenStudy (anonymous):

You just block out the row and the column. So, the 2 is in row one column one so you block out row one and column one. The 3 is in row one column two, so you block out row one and column two. Etc.

OpenStudy (anonymous):

The reason it's +3 instead of -3 is, for the second column you have to subtract instead of adding. So they cancel out.

OpenStudy (anonymous):

thanks